OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7. y = 4x + 10 y = 1over 4x + 5 over 2 y = -4x - 6 y = -1 over 4x + 3 over 2

OpenStudy (hoblos):

parallel lines have same slope the slope of the line (y = 4x + 7) is 4 so the slope of our line is 4 too now apply the formula y-2 = 4(x+2) y-2 = 4x +8 y= 4x +10